OpenStudy (anonymous):

Solve the system, using substitution. One point for correct steps, one point for correct ordered pair. x = 4y and 2x + 3y = 44 Solve the system, using substitution. One point for correct steps, one point for correct ordered pair, (a,b). a + b = 4 and 2a + 3b = 11

OpenStudy (cornzyblack):

Since x = 4y ----- equation 1 2x + 3y = 44 ----- equation 2 Substitute for x in equation 2 that is, 2(4y) + 3y = 44 8y + 3y = 44 11y = 44 y = 44/11 y = 11 substituting y in equation 1 then, x = 4(11) x = 44